OpenStudy (anonymous):

Find the vertex, focus, directrix, and focal width of the parabola. x = 4y^2

OpenStudy (anonymous):

I put Vertex: (0, 0); Focus: ((1/16),0) ; Directrix: x = 1/16 ; Focal width: 16 Which was wrong.

OpenStudy (agent0smith):

Those answers don't help, show your work/attempts.

OpenStudy (anonymous):

looks like your focus is on the ruddy directrix itself. :>

OpenStudy (agent0smith):

\[\large 4p y = x^2\]looks like you got p=1/16 The directrix and focus intersect, which should tell you that there's a problem.